OpenStudy (anonymous):

Graph the two equations to solve the system. What is the solution? 3x + 2y = 6 and y = -3/2 x - 1 A. (2, 0) B. (1, 1.5) C. no solution D. infinitely many solutions

OpenStudy (anonymous):

C. no solution 3x + 2y = 6 solve it for y, we get y=-3/2x+3 -----(A) y = -3/2 x - 1 ------(B) from (A) and (B), both the lines are parallel as they have same slope(-3/2). so their is no solution for this system of equations